# Diverse Perspectives on AI: Examining People’s Acceptability and Reasoning of Possible AI Use Cases

### v1

@@ 4. Analysis Methods
4.1. Multiple Choice Analysis
We analyzed the judgment and confidence ratings by mapping judgment (Q1, Q5) to 1 (“Should be developed”, “Would use”) or -1 (“Should not be developed”, “Would not use”) and confidence (Q2, Q6) to a scale from 1 to 5. We used numerically converted judgment, confidence, and combined (judgmentconfidence; -5 to 5) values as dependent variables in our analysis. We used repeated-measures ANOVAs to understand the differences in mean responses between conditions/groups and linear mixed effects regression models (lmer) to better understand the effects of specific factors. We included a subject-specific random effect when using ANOVA and regression models and added a use-case-specific random effect when applicable. We factorized demographic responses for analysis with the exception of discrimination chronicity, which we aggregated to a numerical value (Kingsley et al., 2024; Michaels et al., 2019). We also converted responses to AI literacy questions to numerical values for analysis.
4.2. Open-response Analysis
To assess the reasoning methods used by the participants, we analyzed the open-text responses on elaborations to their decisions (Q3) and circumstances in which their decisions would switch (Q4) along the following three dimensions: reasoning types (cost-benefit, rule-based, both, unclear), reference to moral foundations (Care, Fairness, Purity, Authority, Loyalty), and switching conditions (Functionality, Usage, Societal Impact). By analyzing reasoning types and moral values reflected in the participants’ justifications, we aim to characterize how participants made their decisions, and by analyzing various factors such as primary concerns and stakeholders, we aim to discover what aspects were salient for the participants in their decisions.
@@ 4.2. Open-response Analysis
We classified participants’ responses to Q3 (elaboration of judgment) and Q4 (conditions for switching decisions) using OpenAI’s o1-mini666o1-mini-2024-09-12. To validate the model’s classification performance, results were compared with a reference set of 100 samples annotated by three independent annotators, comprised of team members and a professional annotator. Initially, each annotator independently assessed the data, and then consensus was reached through discussion to establish a gold standard set. The inter-rater agreement between the gold standard and o1-mini’s annotations was evaluated using Gwet’s AC1 metric, chosen for its robustness with infrequent labels (Wongpakaran et al., 2013). The agreement levels varied, ranging from moderate to almost perfect with no dimension below moderate. Annotations for cost-benefit reasoning, rule-based reasoning, and authority reached near-perfect agreement; purity and usage achieved substantial agreement; the rest showed moderate agreement. Due to a lack of sufficient test samples and minimal occurrences in the annotated data, the moral value dimension Loyalty was excluded from further analysis. The annotations were converted into a one-hot encoding format for statistical analysis. See Appendix B for further details.
5. Findings
@@
5.1. RQ1. Perceptions and Disagreements of Use Cases
5.1.1. Use Case Factors
In our analysis, we investigated the effects of use cases on participants’ judgments of two categories—professional and personal—along with ten specific use cases on judgments and confidence about their development and usage. Figure 2 illustrates significant differences in acceptability depending on the categories. Notably, personal use cases had higher acceptability (; ) than professional use cases (; ). These differences are more pronounced when judgments are weighted by confidence. Although levels of confidence did not differ significantly between categories, they were fairly high for both categories, and personal category exhibited slightly lower confidence for usage.
Professional Use Cases
Exploring specific use cases within the professional category, we observed that Elementary School Teacher AI (; ) had the lowest acceptability for both types of judgments followed by Lawyer AI () and Telemarketer AI (). Interestingly, IT Support Specialist AI had the highest acceptability (; ). While all other use cases showed higher acceptability for usage over development, Telemarketer AI uniquely had higher acceptance for development over usage (; ). Additionally, confidence on using Telemarketer AI () was significantly () higher than that of Government Eligibility Interviewer AI (), which had the lowest confidence in usage. Near 0 mean for development of Lawyer AI () and usage for Telemarketer AI () suggest disagreement within judgments. Moreover, Elementary School Teacher AI uniquely unacceptable across both acceptability judgments (, ) underscoring a unique characteristic perhaps related to care. See Figure 3 for further details
Personal Use Cases
In personal use scenarios, Digital Medical Advice AI (; ), reflecting high risk level, consistently had lower acceptance across judgment types, compared to all other use cases. However, Customized Lifestyle Coach AI had the lowest confidence across both judgments (; ). Nutrition Optimizer (; ) had the highest mean acceptance across the two acceptability judgments. Interestingly, unlike the professional use cases, which had slightly higher acceptance for usage, personal use cases had lower acceptance for usage in general compared to development. Notably, many personal use cases had substantially high acceptance, especially compared to professional use cases.
@@ 5.1.1. Use Case Factors
| Use Case Variation | Judgment | Confidence | Judgment | Confidence |
| Professional | ||||
| Education Level | ||||
| Intercept | ||||
| Personal | ||||
| EU AI Risk Level | ||||
| Intercept | ||||
5.1.2. Use Case Variations
When selecting use cases, we used two underlying variations: entry level of education required for professional use cases and EU AI risk levels for personal use cases. As risk levels and required education increased, we observe consistent negative effects on judgments, with personal use cases (; ) showing stronger effects compared to professional scenarios (), where only development judgments were significantly associated. Again, acceptance for personal use cases (; ) were higher than professional use cases ( ) as conveyed by the intercepts. Confidence ratings remained consistently high across all conditions (intercepts ¿ 3.97), though they showed a small but significant decrease with increasing risk levels in personal use cases (; ), while professional use cases showed no significant impact on confidence. Thus, while education level required for entry does show some significant effect, the EU AI risk level had consistent effects with higher significance.
5.1.3. Disagreements
Observing the standard deviation and visualization of the distribution shows further understanding of possible disagreements among use cases. We compare the judgments weighted by confidence to understand not only the differences in judgment but also their strength. Interestingly, the use cases with four highest disagreements in both judgments were all professional uses in order of Telemarketer (; ), Elementary School Teacher (; ), Lawyer (; ), and Government Eligibility Interviewer AI (; ). These four use cases were followed by Digital Medical Advice AI (, ). The use cases with the lowest disagreements were surprisingly Nutrition Optimizer (; ) followed by IT Support Specialist AI (; ). These results underscore the general controversy of professional usages but also shows that acceptability is highly use case dependent.
5.2. RQ2. Impact of Personal Factors on Acceptability Judgment
5.2.1. Demographic Factors
@@
In analyzing the demographic factors influencing judgments on the development and usage of AI use cases, several significant trends emerged from the data. First, we observed that the demographic variables had less than 0.5 correlation, except for age 65+ and Retired employment status. Across both categories of use cases, age was positively associated with confidence in acceptability judgment of usage for age 25-34 (, ) and 55-64 (, ). Race also had notable influences; specifically, Asian participants exhibited significantly lower confidence in both development and usage judgments (, ; , ), particularly in professional contexts.
Gender emerged as a crucial determinant, with non-male participants consistently showing negative judgments across both development (, ) and usage (, ), indicating potential discrepancies in perception or experience with AI applications. Liberal views, especially among those identifying as strongly liberal, were associated with negative judgments across both categories of use cases (, ; , ), suggesting a skeptical stance towards AI’s prevalence and role. Employment hours also contributed, with individuals working 40+ hours per week displaying a positive association with development judgments (, ), suggesting more exposure or reliance on AI use cases. High experience of discrimination chronicity was significantly related to lower acceptance of development, , . These findings highlight the significant role of demographic factors in shaping perceptions and attitudes toward AI technologies. See Table 23 in the Appendix for ANOVA results.
@@ 5.2.2. AI Literacy
We identified a correlation greater than 0.5 among three AI literacy aspects: awareness, usage, and evaluation. These were aggregated into a single factor, AI Skills. As shown in Table 4, understanding of AI Ethics was associated with lower acceptability for both personal (, ; , ) and professional (). However, across both categories of use cases, high Generative AI Usage Frequency resulted in higher acceptance for both professional (, ; , ) and usage acceptance (, ; , ). Notably, for personal use cases, AI Skills was positively associated with confidence of judgments (, for both development and usage) and judgment weighted by confidence (, ), while Generative AI Limitation Familiarity was positively associated with confidence for professional usage (, ; , for usage). These results implicate that different understandings of and experiences with AI can impact judgments of acceptability, corroborating previous findings (Kramer et al., 2018).
5.3. RQ3. Factors in Participant Rationale
@@
To provide deeper insights to our analyses of acceptability judgments, we also examined open-text rationales (Q3) for judgments of development (Q1) and conditions for switching their decisions (Q4).
5.3.1. Decision-making Types
@@
As we defined in § 4.2, we focus on two distinct reasoning types for decision-making, distinguished by consideration of outcome versus consideration of value inherent in action: cost-benefit reasoning (e.g., “it gives more people access to medical advice and treatment”, P365) and rule-based reasoning (e.g., should not be developed because “human interaction is better”, P249). As shown in Figure 5, generally, participants used more cost-benefit reasoning with highest percentage for IT Support Specialist (91.0%) and Nutrition Optimizer (92.8%). This result is particularly interesting as we observed these two use cases to have the lowest disagreement (see § 5.1.1) and signifies that perhaps unified reasoning type leads to less disagreement.
On the other hand, rationales for Elementary School Teacher AI contained most percentage of rule-based reasoning (30.0%) followed by Lawyer AI (22.0%). Interestingly, Elementary School Teacher AI and Lawyer AI were the use cases with the lowest development acceptability. These results suggest that less acceptable AI use cases might trigger more rule-based reasoning, which could, however, be due to the lack of consensus and rules around AI, especially those that are positive. this suggests the use of rule-based reasoning might correlate with judging AI use cases to be more unacceptable. These results suggest that the use of rule-based reasoning might correlate with judging AI use cases to be more unacceptable.
5.3.2. Moral Foundations
@@
Beyond reasoning types, we explored moral foundations to provide insights into what values are relevant for AI use case decisions (Figure 6). For example, P12 responded that Elementary School Teacher AI use case should be developed because it “could give elementary schooling to children who are bed ridden…”, which was annotated with both values of Care (focusing on the well-being and nurturing of bed-ridden children) and Fairness (focusing on fair access to education). Upon analysis, we observe that Care (i.e., dislike of pain of others, feelings of empathy and compassion toward others) was the most prevalent moral value in participants’ rationales across the use cases. Of note, Care could impact the acceptability of AI use cases in both directions, as conveyed by P385 who noted that Customized Lifestyle Coach AI should be developed because “it may help improve some people’s health” but would change their decision if “it caused harm to even one person.” While our choice of medical domain in personal usage could have impacted the distribution, care was still the most prevalent considering only professional use cases.
@@
Participants’ rationales evoked fairness value more than care for two use cases, Lawyer AI (0.62) and Government Eligibility Interviewer (0.74) respectively. These results could be due to the characteristics of the use cases, such as their main purpose and function, as noted by P88, Government Eligibility Interviewer should be developed because “it might be less biased and therefore more fair in it decisions (sic)”. Authority was most apparent in participant rationales for Lawyer AI (0.37) and Purity for Flavorful Swaps (0.35).
5.3.3. Switching Conditions
@@
We have thus far explored firmness of the decision through levels of confidence provided by users. We further explore the flexibility of participants’ judgments to understand possible mitigation of disagreements through concerns expressed in conditions under which they would switch their decisions (Figure 7). Overall, Functionality (53.3%; e.g., Medical Adivce AI should not be developed if it “consistently or had a high percentage of failure to diagnose correctly.”, P373) was the most commonly noted concern that would switch participants decision. This was followed by Societal Impact (42.3%; e.g., Lawyer AI should not be developed if “it puts too many human lawyers out of work”, P97), Usage (31.8%; e.g., Government Eligibility Interviewer should be developed if “it was only used to read and screen applications but not for making decisions”), Not Applicable (0.02%; e.g., will not change decision).
Interestingly, Societal Impact (50.6%) was more prevalent followed by Functionality (46.4%) for professional use cases whereas Functionality (54.4%) was the most frequently mentioned concern for personal use cases. One of the reasons for frequent mention of Societal Impact in professional use cases could be due to concerns of labor replacement for professional use cases: as described by P296, if Elementary School Teacher AI “was to replace teachers with the ai to save money”, they would switch decision from should be developed to should not be developed. Moreover, we observe that Societal Impact was the most prevalent concern for those (31%) who thought the use case was unacceptable across all use cases and was especially prevalent for Elementary School Teacher AI (0.47). These diverging results show the importance of understanding granular concerns by use cases to effectively address the relevant issues. Participants’ responses that were identified as being both rule-based and concerned about societal impact expressed rules such as necessity of humanness as expressed by P22, “a human touch is 100% necessary” when discussing Elementary School AI, whether it be due to belief that “AI would lack empathy” (P5) or that “humans need human interactions” (P44).
| DEV ( (SE)) | USAGE ( (SE)) | |||||
| Rationale Factors | Judg. | Conf. | Judg.Conf. | Judg. | Conf. | Judg.Conf. |
| Reasoning Type | ||||||
| Cost-benefit | ||||||
| Rule-based | ||||||
| Moral Value | ||||||
| Care | ||||||
| Fairness | ||||||
| Purity | ||||||
| Authority | ||||||
| Switching Condition | ||||||
| Functionality | ||||||
| Usage | ||||||
| Societal Impact | ||||||
5.3.4. Rationale Factors and Judgment
The results of analyzing the influence of rationale factors on judgments of acceptability indicated several significant effects (Table 5). We first verified that the factors had less than 0.5 correlation with each other. For acceptability of development, use of Cost-benefit reasoning in rationale was associated with increased acceptability (), while Rule-based reasoning was negatively associated with acceptability (). Interestingly, for usage, Cost-benefit reasoning was negatively associated with acceptability () but positively for judgment weighted by confidence (). Rule-based reasoning indicated lower usage acceptability for judgment weighted by confidence, consistent with acceptability of development. In the moral value category, Fairness in rationales had positive associations with acceptance (, ; judgment weighted by confidence).
Analyzing switching conditions showed that participants who answered negatively to the development of use cases mentioned the concerns about Functionality () and Societal Impact to switch their decisions (). However, those who were positive towards development of use case indicated emphasis on Usage () as a condition to reverse their decisions. Thus, our findings highlight diverse perspectives on requirements and concerns of AI use cases, especially with varying perception of acceptability.
6. Conclusion and Discussion

### v2

@@ 4 Acceptability & Reasoning
4.1 Multiple Choice Analysis
We analyzed the judgment and confidence ratings by mapping judgment (Q1, Q5) to 1 (“Should be developed”, “Would use”) or -1 (“Should not be developed”, “Would not use”) and confidence (Q2, Q6) to a scale from 1 to 5. We used numerically converted judgment, confidence, and combined (judgmentconfidence; -5 to 5) values as dependent variables in our analysis. We used repeated-measures ANOVAs to understand the differences in mean responses between conditions/groups and linear mixed effects regression models (lmer) to better understand the effects of specific factors. We included a subject-specific random effect when using ANOVA and regression models and added a use-case-specific random effect when applicable. We factorized demographic responses for analysis with the exception of discrimination chronicity, which we aggregated to a numerical value (Kingsley et al. 2024; Michaels et al. 2019). We also converted responses to AI literacy questions to numerical values for analysis.
4.2 Open-response Analysis
Background: Moral Decision Making
To understand decision-making in AI use cases, we draw on moral psychology and dual system theory. We examine two decision-making systems: cost-benefit reasoning, which assesses outcomes and consequences, and rule-based reasoning, focusing on norms, rules, and virtues (Cushman 2013; Cheung, Maier, and Lieder 2024). These correspond to utilitarian reasoning (maximizing good) and deontological reasoning (adhering to moral duties and rights), respectively. Additionally, we apply moral foundations theory (Graham et al. 2008) to identify values and potential moral conflicts in AI development.
To assess the reasoning methods used by the participants, we analyzed the open-text responses on elaborations to their decisions (Q3) and circumstances in which their decisions would switch (Q4) along the following three dimensions: reasoning types (cost-benefit, rule-based, both, unclear), reference to moral foundations666We used the five foundational dimensions: Care, Fairness, Loyalty, Authority, and Purity. Although these dimensions have been updated to encompass a broader range of values beyond WEIRD (White, Educated, Industrialized, Rich, and Democratic) populations (Atari et al. 2023), we selected this version for survey brevity. (Care, Fairness, Purity, Authority, Loyalty), and switching conditions (Functionality, Usage, Societal Impact). By analyzing reasoning types and moral values reflected in the participants’ justifications, we aim to characterize how participants made their decisions, and by analyzing various factors such as primary concerns in switching condition, we aim to discover what aspects were salient for the participants in their decisions.
@@ 4.2 Open-response Analysis
We classified participants’ responses to Q3 (elaboration of judgment) and Q4 (conditions for switching decisions), totaling 985 samples for each question, using OpenAI’s gpt-4o777gpt-4o-2024-11-20. To validate the model’s classification performance, results were compared with a reference set of 100 samples annotated by three independent annotators, comprised by members of the research team and a professional annotator. Initially, each annotator independently assessed the data, and then consensus was reached through discussion to establish a gold standard set. The inter-rater agreement between the gold standard and o1-mini’s annotations was evaluated using Gwet’s AC1 metric, chosen for its robustness with infrequent labels (Wongpakaran et al. 2013). While the agreement levels varied, ranging from almost perfect to moderate (0.98–0.57), all dimensions had above substantial agreement except Societal Impact. Annotations for cost-benefit reasoning, rule-based reasoning, and authority reached near-perfect agreement. Due to minimal occurrences in both human and LLM annotations, the moral foundation dimension Loyalty was excluded from further analysis. The annotations were conducted based on presence or absence of the values and were converted into binary format for statistical analysis. See Appendix C for further details on agreement and automatic annotation settings.
5 Findings
@@
5.1 RQ1. Use Case Perceptions & Disagreements
In our analysis, we investigated the effects of use cases on participants’ judgments using our ten use case vignettes. Overall, acceptability statistically differed among the two categories (; ). Notably, personal health use cases had higher acceptability (; ) than labor-replacement use cases (; ). See Figure 7 in Appendix B for additional category comparison results.
Labor-replacement Use Cases
Exploring specific use cases within the labor-replacement category (Figure 2), we observed that Elementary School Teacher AI (; ) had the lowest acceptability for both types of judgments followed by Lawyer AI () and Telemarketer AI () where their near zero mean suggest disagreement within judgments. Interestingly, IT Support Specialist AI had the highest acceptability (; ) despite human replacement potential and median education level.
Personal Health Use Cases
In personal health use scenarios, Digital Medical Advice AI (; ), reflecting high risk usage, consistently had lower acceptance across judgment types, compared to all other use cases. Nutrition Optimizer (; ) had the highest mean acceptance across both acceptability judgments. Interestingly, unlike the labor-replacement use cases which had slightly higher acceptance for usage, personal use cases had lower acceptance for usage in general compared to development.
@@
| Judg. | Conf. | Judg. | Conf. | |
| Labor-replacement | ||||
| Coeff. | ||||
| Personal Health | ||||
| Coeff. | ||||
Use Case Variations
When selecting use cases, we used two underlying variations: entry level of education required for labor-replacement use cases and EU AI risk levels for personal health. As risk levels and required education increased, we observe consistent negative effects on judgments, with personal health use cases showing stronger effects (; ) compared to labor-replacement scenarios (), where only development judgments were significantly associated. Confidence ratings showed a small but significant decrease with increasing risk levels in personal health use cases (; ), while labor-replacement use cases showed no significant impact on confidence.
Disagreements
We compare the standard deviation of judgments weighted by confidence to understand possible disagreements and their strength among use cases. Interestingly, the use cases with four highest disagreements in both judgments were all labor-replacement uses in order of Telemarketer (; ), Elementary School Teacher (; ), Lawyer (; ), and Government Eligibility Interviewer AI (; ). These four use cases were followed by Digital Medical Advice AI (, ). The use cases with the lowest disagreements were surprisingly Nutrition Optimizer (; ) followed by IT Support Specialist AI (; ).
RQ1 Takeaways
Our results show that there are significant variation in judgments based on use case characteristics such as category, EU-defined risk level, and, to some extent, required education level for labor-replacement cases. The uniquely negative response to the Elementary School Teacher AI highlights potential concerns specific to care work. High disagreement in labor-replacement scenarios underscores the need for cautious integration of AI into existing roles. In contrast, the consistent positive judgments for IT Support Specialist AI suggest that not all labor-replacement use cases are viewed equally, indicating the need for nuanced understandings of acceptability. We further examine this variability in § 5.3.
@@
5.2 RQ2. Impact of Personal Factors on Acceptability Judgment
Demographic Factors
As shown in Table 3, several demographic factors significantly influenced use case judgments.888Prior to analysis, we observed that the independent variables in the model had less than 0.5 correlation, except for age 65+ and Retired employment status. Across both categories of use cases, certain age groups were positively associated with confidence in usage: 25-34 (, ) and 55-64 (, ). Race also had notable influences; specifically, Asian participants exhibited significantly lower confidence in both development and usage judgments (, ; , ), particularly in labor-replacement contexts.
Gender emerged as a crucial determinant, with non-male participants consistently showing negative judgments (, ; , ), indicating potential discrepancies in perception or experience with AI applications. Liberal views, especially among those identifying as strongly liberal, were associated with negative judgments across both categories of use cases (, ; , ), suggesting a skeptical stance towards AI’s prevalence and role. Employment hours also contributed, with individuals working 40+ hours per week displaying a positive association with development (, ), suggesting more exposure or reliance on AI use cases. High experience of discrimination chronicity was significantly related to lower acceptance of development (, ). See Table 25 in the Appendix for ANOVA results.
@@
We identified a correlation greater than 0.5 among three AI literacy aspects: awareness, usage, and evaluation. To avoid multicollinearity, we aggregated them into a single factor, AI Skills. As shown in Table 4, understanding of AI Ethics was associated with lower acceptability for both personal health (, ; , ) and labor-replacement (). However, across both categories, high Generative AI Usage Frequency resulted in higher acceptance (Labor-replacement – , ; , , Personal Health – , ; , ). Notably, for personal health use cases, AI Skills was positively associated with confidence of judgments (, ), while Generative AI Limitation Familiarity was positively associated with confidence for labor-replacement usage (, ; , ).
RQ2 Takeaways
Our findings highlight the significant role of lived experiences and backgrounds, such as age, race, gender, political view, employment, and discrimination experience, in shaping perceptions and attitudes toward AI technologies. Moreover, our results indicate that different understandings of and experiences with AI can impact judgments of acceptability, corroborating previous findings (Kramer et al. 2018).
| DEV ( (SE)) | USAGE ( (SE)) | ||||||
| Judg. | Conf. | Judg.Conf. | Judg. | Conf. | Judg.Conf. | ||
| Reasoning Type | |||||||
| Cost-benefit | |||||||
| Rule-based | |||||||
| Moral Value | |||||||
| Care | |||||||
| Fairness | |||||||
| Authority | |||||||
| Switching Condition | |||||||
| Usage | |||||||
| Societal Impact | |||||||
5.3 RQ3. Factors in Participant Rationale
@@
To deepen our analysis of use case acceptability, we examined open-text rationales (Q3) and decision-switching conditions (Q4) related to development judgments (Q1).
Decision-making Types
@@
As we defined in § 4.2, we focus on two distinct reasoning types for decision-making: cost-benefit reasoning, which emphasizes outcomes (e.g., “it gives more people access to medical advice and treatment”, P365), and rule-based reasoning, which reflects values inherent in the action itself (e.g., should not be developed because “human interaction is better”, P249). As shown in Figure 4, generally, participants used more cost-benefit reasoning, especially for the IT Support Specialist (91.0%) and Nutrition Optimizer (92.8%) use cases. This result is particularly interesting as we observed these two use cases to have the lowest disagreement (see § 5.1), suggesting unified reasoning type may lead to more consistent judgments. On the other hand, rationales for Elementary School Teacher AI contained most percentage of rule-based reasoning (30.0%) followed by Lawyer AI (22.0%), both of which were the two use cases with lowest development acceptability.
Through further analysis of the influence of rationale factors on judgments using a mixed-effects model (Table 5), we found that use of cost-benefit reasoning in rationale is positively associated with development acceptability (), while rule-based reasoning is negatively associated (). Interestingly, for usage, cost-benefit reasoning was negatively associated with acceptability () but positively for judgment weighted by confidence (). This suggests that, although cost-benefit considerations may reduce usage acceptance, they boost confidence when judgments are favorable.
The significant negative association of rule-based reasoning with judgments indicate that for certain contexts, the inherent action of using AI is viewed negatively. Participants cited diverse concerns related to AI itself (e.g., “because it would not have human sympathy”, P16), human needs (e.g., “Humans need human interactions in order to learn properly”, P44), societal impact (e.g., “Overreliance on AI …”, P132), and morality (e.g., “Having artificial intelligence try and sell you things is immoral”, P13). These results suggest that AI’s acceptability heavily lies in its positive outcomes but can be outweighed by established rules.
Moral Foundations
@@
Beyond reasoning types, we explored moral foundations to provide insights into what values are relevant for AI use case decisions (Figure 5). For example, P12 responded that Elementary School Teacher AI use case should be developed because it “could give elementary schooling to children who are bed ridden…”, which was annotated with both values of Care (focusing on the well-being and nurturing of bed-ridden children) and Fairness (focusing on fair access to education). Upon analysis, we observe that Care (i.e., dislike of pain of others, feelings of empathy and compassion toward others) was the most prevalent moral foundation in participants’ rationales across the use cases in both categories (48%). Interestingly, Care could be invoked in both positive and negative regards for AI, as conveyed by P385 who noted that Customized Lifestyle Coach AI should be developed because “it may help improve some people’s health” but would change their decision if “it caused harm to even one person.”
@@
Although Care was the most dominant moral foundation overall, Fairness emerged more prominently in context-specific evaluations such as Lawyer AI (41%) and Government Eligibility Interviewer AI (66%). These results could be due to the characteristics of the use cases, such as their main purpose and function; as noted by P88, Government Eligibility Interviewer should be developed because “it might be less biased and therefore more fair in it decisions (sic)”. Authority was most apparent in participant rationales for Lawyer AI (28%) and Purity for Flavorful Swaps (27%). Moreover, Fairness in rationales had positive associations with acceptance (, ; judgment weighted by confidence; see Table 5).
Switching Conditions
@@
We further explored the flexibility of participants’ judgments to understand possible mitigation of disagreements through criteria for switching their decisions (Figure 6). Functionality (53%; e.g., Medical Adivce AI should not be developed if it “consistently or had a high percentage of failure to diagnose correctly.”, P373) was the most commonly noted condition for switching decisions in both directions (positive to negative and vice versa). This was followed by Usage (40%; e.g., Government Eligibility Interviewer should be developed if “it was only used to read and screen applications but not for making decisions”), Societal Impact (36%; e.g., Lawyer AI should not be developed if “it puts too many human lawyers out of work”, P97), and Not Applicable (7%; e.g., will not change decision).
Interestingly, for labor-replacement use cases Societal Impact (45%) was more frequently mentioned as switching conditions followed by Functionality (43%), whereas Functionality (53%) was more frequently mentioned than Societal Impact (37%) for personal use cases. Frequency of Societal Impact in labor-replacement use cases could be closely linked to concerns of labor replacement: as described by P296, if Elementary School Teacher AI “was to replace teachers with the ai to save money”, they would switch their decision from positive to negative. Moreover, for all use cases except Elementary School Teacher AI, participants tended to switch from positive to negative decisions for lack of functionality reasons. In contrast, for Elementary School Teacher AI, the most common shift towards acceptability when the use case showed a positive societal impact (38%). As shown in Table 5, mentions of Societal Impact as conditions to switch decisions were more negatively associated with judgments (, ; judgment confidence). However, emphasis on Usage () as a condition to reverse their decisions was positively associated.
RQ3 Takeaways
Reasoning types varied by context, with rule-based reasoning more common in contested use cases and negatively associated with acceptability, while cost-benefit reasoning showed a positive association. The moral foundation of Care was especially salient, highlighting its importance in AI judgments. When explaining what might change their decisions, participants most frequently cited Functionality for personal health use cases and Societal Impact for labor-replacement, underscoring the context-dependent nature of these concerns, especially to be considered when mitigating disagreements.
6 Conclusion and Discussion